OpenStudy (anonymous):

Solve 4x + 3y = 13 x + 5y = 33 a. (2, 7) b. (–2, 7) c. (2, –7) d. (–2, –7)

OpenStudy (cwrw238):

first multiply 2nd equation by -4 and add to eliminate x - not forgetting to multiply every term by -4

OpenStudy (anonymous):

4x + 3y = 13 -4(x + 5y = 33) 4x + 3y = 13 -4x - 20y = -132 -17y = -119 y = 7 4x + 3(7) = 13 4x + 21 = 13 - 21 -21 4x = -8 x = -2

OpenStudy (anonymous):

@cwrw238 did i do it right?

OpenStudy (cwrw238):

mistake in seventh line should be 4x+ 5(7)

OpenStudy (cwrw238):

sorry - im wrong - you got it right

OpenStudy (cwrw238):

answer is b)